4 x +6 x =9 x . Find ‘x’.

4x+6x=9x. Find ‘x’.

Grade:Select Grade

3 Answers

Rajorshi Koyal
8 years ago
0.Approv
Chandrabhushan Reddy Chigarapalli
7 years ago
Divide both sides with 6x.You will get      
(2/3)x +1 =(3/2)x
Assume (2/3)x =y.
By solving you will get a quadratic in variable y.
By using quadratic formula you will be able to get the value of x.
Shreyansh Pampaniya
4 years ago
4^x + 6^x = 9^x
1 + (3/2)^x =(3/2)^2x
 Let (3/2)^x be 't'
1 + t = t^2
t^2 - t -1 =0
 
t=[ 1+- root (5) ] / 2
 (3/2)^x = t
 x = lnt/ln(3/2)    [substitute value of t]
 
 (3/2)